$ 2,700 3.270 Supplies. Prepaid insurance Accumulated depreciation Equipment Long-term notes payable. 1,500 6,400 . . $ 1,080 845 7,000 3,000 5,205 2,900 Service fees eamed... . 62,400 15,000 33,400 4,700 3,250 720 $79,530 $79,530

Prepare a detailed Balance Sheet and also calculate the following ratios:

A) Current Ratio

B) Working Capital

C) Assets Turnover

D) Fixed Assets Turnover

E) A/R Turnover

F) Debt Ratio
Answer #1
A B C D E F G
2
3 A)
4
5 Current Ratio =Current Assets / Current Liabilities
6
7 Cash $2,700
8 Accounts Receivable $3,270
9 Supplies $5,060
10 Prepaid insurance $1,500
11 Current Assets $12,530 =SUM(D7:D10)
12
13 Accounts Payable $845
14 Current Liabilities $845
15
16 Current Ratio =Current Assets / Current Liabilities
17 14.83 =D11/D14
18
19 Hence current ratio is 14.83
20
21 B)
22
23 Working capital =Current Assets - Current Liabilities
24 =12530-845
25 $11,685
26
27 Hence working capital is $11,685
28
29 C)
30
31 Asset Turnover =Total Sales / Total Assets
32
33 Total Sales $62,400
34
35 Current Assets $12,530
36 Equipment $6,400
37 Accumulated depreciation - Equipment ($1,080)
38 Total Assets $17,850 =SUM(D35:D37)
39
40 Asset Turnover =Total Sales / Total Assets
41 $3.50 =D33/D38
42
43 Hence Asset Turnover is $3.50
44
45 D)
46
47 Fixed asset turnover =Total Sales/ Net fixed assets
48
49 Equipment $6,400
50 Accumulated depreciation - Equipment ($1,080)
51 Net Fixed Assets $5,320 =SUM(D48:D50)
52
53 Revenue $62,400
54
55 Fixed asset turnover =Total Sales/ Net fixed assets
56 11.73 =D53/D51
57
58 Hence Fixed asset turnover is 11.73
59
60 E)
61
62 A/R Turnover =Credit Sales / Accounts Receivable
63 =62400/3270
64 19.08
65
66 Hence A/R Turnover 19.08
67
68 F)
69
70 Debt Ratio =Total Debt / Total Assets
71
72 Total Debt =Long term notes payables
73 $7,000
74
75 Total Assets $17,850
76
77 Debt Ratio =Total Debt / Total Assets
78 0.39 =D73/D75
79
80 Hence Debt Ratio is 0.39
81